Two electric dipoles of moment P and 64P placed in opposite direction on a line at a distance of 25cm the electric field will be zero at point between the dipole whose distance from the dipole of moment P is

Grade:12th pass

1 Answers

Vaishali
5 years ago
Suppose a point N (neutral point) lies between the two dipoles (p and 64 p) which is at x distance from dipole moment pHence distance from dipole moment 64p= 25-xAt NE.F. due to dipole p = E.F. due to dipole 64p2kp/x^3=2k(64p)/(25-x)^31/x^3= 64/(25-x)^3x^3= (25-x)^3/64x= [(25-x)^3/64]^1/3x= 25-x/44x+x=25x=5cm